Org Pvt. Ltd. is considering two mutually exclusive capital investments. The project’s

expected net cash flows are as follows:

Expected Cash Flows

Year Project A Project B

0 -400 -575

1 95 150


2 110 200

3 118 250

4 125 275

5 140 230

6 150 180

a. If you were told that each project’s cost of capital was 10%, which project should be

selected using the NPV criteria?

b. What is each project’s IRR?

c. What is the regular payback period for these two projects?

d. What is the profitability index for each project if the cost of capital is 12%?


1
Expert's answer
2021-05-27T16:17:50-0400

a. "NPVa = -400+95\/1.1+110\/1.1^2+118\/1.1^3+125\/1.1^4+140\/1.1^5+150\/1.1^6 = 122.9."


"NPVb = -575+150\/1.1+200\/1.1^2+250\/1.1^3+275\/1.1^4+230\/1.1^5+180\/1.1^6 = 346.73."


b. IRR is a cost of capital, at which NPV = 0. IRR is higher for project B.


c. Payback period is shorter for B.


d. Profitability index is higher for B.